- Estimate energy (in joule) of one human heartbeat.

-Estimate energy (in joule) of a hummingbird in flight.

-How many joule are in 1 kilowatt-hour?

-How many joule are in 1 Food Calorie?
Answer #1

Answer:

(a) Average power output of the human heart is 1.33 W. We can convert power into energy, watt (W) = joule/sec

Therefore, 1.33 W = 1.33 J/s.

Hence, a human heart consumes 1.33 J of energy per second and for one minute it consumes 60 x 1.33 J = 79.8 J.

Now, average heartbeat rate of a healthy human is 72 per minute. Then, the heart beats 72 times in a minute and consumes 79.8 J of energy. Therefore, 79.2 J/72 = 1.108 J energy per one heartbeat.

(b) Taking the average mass of a humming bird is m = 4 g or 0.004 kg and its speed is 22 m/s or 79 km/h.

Therefore, the kinetic energy of the humming bird in flight is k = 1/2 mv2 = 1/2 (0.004 kg) (22 m/s)2 = 0.968 J.

(c) 1 watt = 1 Joule/s so one kilo watt is equal to 1000 times of watt then 1kW = 1 kJ/s. Now,

1 kWh = (1kJ/s) h = (1 kJ/s) (3600 s) = 3600 kJ or 3600000 J or 3.6 x 106 J.

(d) One calorie of food has 4.184 J of energy